OpenStudy (anonymous):

write this equation in slope intercept form: -y-5x=25

OpenStudy (anonymous):

if you answer this i will give you a medal!

OpenStudy (anonymous):

y = -5x - 25

OpenStudy (anonymous):

-y-5x=25 +5x +5x -y=5x+25 /-1 /-1 y=-5x-25 <------------- TADA

OpenStudy (anonymous):

do you have to show your work on the hw u have???

OpenStudy (anonymous):

if u dont i can just give u the answer

OpenStudy (anonymous):

i don't have to show work, i love my teacher. all i need is the answer thanks!

OpenStudy (anonymous):

sure, i will just give the answer next time

OpenStudy (anonymous):

ok sounds great!